How many liters of a mixture containing 80% alcohol should be added to 5 liters of 20% alcohol to yeild a 30% alcohol solution?
Answer by lwsshak3(11628)

How many liters of a mixture containing 80% alcohol should be added to 5 liters of 20% alcohol to yeild a 30% alcohol solution?
**
let x = liters of 80% alcohol to be added
5+x = liters of 30% solution
..
80%x+20%*5=30%(5+x)
.8x+1=1.5+.3x
.5x=.5
x=1
ans:
liters of 80% alcohol to be added = 1